ALERT REPORT 2: Week 2.  Sunday January 9, 2005 - Saturday January 15, 2005
Prepared: January 15, 2005
 

THINGS WE NEED:
*The items will be left on the list until they are physically received in Alert.*


Please note:  Inventories are current as of the date the report is written.


AEROSOL PROGRAM


Aethalometer 1
Last blank on day 003j, 2005.
O.K.

Aethalometer 2
Currently not in use.

PSAP Program
O.K.

Nephalometer
O.K.

CNC Counter
O.K.

High Volume Samplers
O.K.
Last Blank: 355j, 2004
0 filter(s) in inventory


Toxics Sampler 
O.K.
6 PUF tin(s) + 3 Blank(s) remaining in inventory
3 Toxics case(s) remaining in inventory
Last blank: 362j, 2004. 


MET PROGRAM

Ambient Temperature Sensor [Campbell 107F]
O.K.

Stevenson Screen and Dew Cells
Changed monitoring TLD for Health Canada on 014j at 1601z.
O.K.

Tall Tower 10m and Lab Platform 10m RM Young Wind Sensors 
O.K.


GAS PROGRAM

NDIR
O.K.
NDIR Moisture Trap Sample
Last trap change: day 355j, 2004.

CAPMoN Ozone Teco #49C-73452-373
O.K.

CAPMoN Ozone Calibration Unit
O.K.

Old PAN GC
Found PAN integrator out of paper upon arrival to lab on 012j. Re-programmed integrator and running at 1427z.
O.K.

New PAN GC
New PAN GC, Cal/Valve unit, and laptop shipped south for service on 349j.
O.K.

CH4/CO2/N2O/SF6 GC
GC1 sequence offline on 014j from 1530z - 1542z for rotor change.
O.K.

CO GC
O.K.

CFC 11, 12, 13 GC
O.K.


Support Gas Tank Inventory
- current as of 010j, 2004
	
	Nitrogen: 2 tank(s)
	Argon: 12 tank(s) (6 large steel tanks; 6 small aluminum tanks)
	P-5: 2 tank(s)
	Zero Air: 7 tank(s)
	Ultra Zero Air: 1 tank(s)
	Hydrogen: 9 tank(s)
	He: 4 tank(s) (5 small aluminum tank)
	CO2 tanks: 0 NDIR tanks
	   

MERCURY PROGRAM

Ambient air (007):
007 offline on 010j from 1847z - 1854z for Ar tank change:
    Old tank off at 1849z (AES722; 150 psi)
    New tank on at 1852z (AES713; 2000 psi)
O.K.

Ambient air (012):
012 offline since 292j.
O.K.

RGM Unit, 1130 0016, 2537A 040:
040 offline on 010j from 1847z - 1913z for Ar tank change and RGM glassware and filter change:
    Old tank off at 1849z (AES722; 150 psi)
    New tank on at 1852z (AES713; 2000 psi)
    Denuder, pyrolyzer, particulate heating and desorb manually on at 1913z
    Denuder, pyrolyzer, particulate heating off and denuder and system cooling on at 1941z.
    System back to Auto at 1952z.
O.K.

RDMQ
2005 QC for 007 and 012 complete up to 013j

Mercury Snow Sampling Experiment
Last samples collected on 181j.
Last blank on 176j.

Particulate Mercury Samplers
Shut down for season.


OTHER PROGRAMS

Health Canada Precipitation Collector
Changed: 003j.
Inventory: 25 month(s)

Health Canada Thermo-Luminescence Detector
Changed: 014j.

Brewers
019:
O.K.
029:
O.K.



INTERNATIONAL PROGRAMS

Weekly 14 CO2 Sample for U. of Heidelberg
Last Sample: 010j
O.K.
Inventory: 11 bottle(s), 22 week(s)

Radon Monitor
Found Radon progrom stopped on 015j at 1815z. Typed radlog log and program up and running at 1816z.
O.K.
Last filter change: Day 003j.

DOAS System
Instrument arrived back on 012j. Will install DOAS during week 3.
O.K.

IGLOOLIK Sampler
HiVol would not restart on 306j. HiVol back up running on 308j at 1520z after fuse change. HiVol 2 not currently able to run. 3 motors sent south for repairs on week 48 and 50.

SEARCH
O.K.


FLASK PROGRAM


Sampling Conditions:  -24.9 C, 0 m/s, R.H. 43.6%, 990.8 mbar.		
Flasks sampled:	12/01/2005	(12j)

Scripps 5L glass greased stopcock, evacuated:  M5037, M5038	
Inventory	4 box(es), 4 week(s)

MSC Ocean Chemistry (Victoria), I.O.S.:  NA	
Inventory	28 flask(s), 28 week(s)

Japanese Halocarbon, steel 5L:  1145, 1310	
Inventory	2 box(es), 4 week(s)

NOAA 2.5L 007 pressurized:  0, 0, 0, 0	
Inventory	18 box(es), 18 week(s)

MSC 2L glass double valve pressurized:  P2-105, P2-106	
Inventory	1 box(es), 2 week(s)

NOAA stainless steel 850mL/2.5L pressurized:  5137L, 5138R	
Inventory	5 box(es), 5 week(s)

York University VOC steel, pressurized:  AES073	
Inventory	6 flask(s), 6 week(s)

CSIRO/MSC 2L glass, pressurized:  P2-236, P2-237	
Inventory	3 box(es), 6 week(s)

SCRIPPS 3 X 5L Oxygen:  199, 316, 524	
Inventory	1 box(es), 2 week(s)

PRL VOC, pressurized:  AL014	
Inventory	9 flask(s), 9 week(s)

U of Heidelberg1L glass pressurized:  22-Q114, 22-Q115	
Inventory	24 flask(s), 12 week(s)

Max Planck Institute - BGC:  J0866, J0867,J0868	
Inventory	3 case(s), 3 week(s)





COMPUTERS



TRANSPORTATION

Fuel Consumption: 135L in December.

Ford (Nuuk)
Tie rod broken on 344j. Mechanics providing a list of parts that will be required to fix Ford and power-steering assist is suggested to prevent this problem occurring again. Truck is currently considered non-serviceable. Parts will be sent up within January.
O.K.

GMC (EC Rider)
Heater now working. Awaiting parts from down south.
O.K.


BUILDINGS
O.K.


SHIPPING

Shipping Received:
NONE


Shipping Sent:
12/01/05
1 case Victoria


MAIL
   
Mail Received:
12/01/05
1 box NOAA SS
2 boxes Scripps 3x5L
4 cases MPI-BGC
1 case DOAS
1 box spare moisture trap
1 parcel Health Canada Thermo-Luminescence Detector
1 parcel mailbags
1 box MPI-BGC spare parts
1 box digital camera


Mail Sent:
1 box Scripps 5L
1 case MSC Orange
1 case MPI-BGC
1 case Toxics
1 box empty PUF Tins
1 box NOAA SS
1 case Hg denuders
1 box pump (Armand)
   

Alert Weekly News

Hello everyone,
 
 It's been a fun week with a few new Vimy members moving in to the house and a few extra people from Trenton that have brought station numbers up a bit and have gotten everyone in a better mood. Vimy is looking to renovate and knock out a few walls (hopefully that will get done while I'm away!) to expand the common room and kitchen, so we're having a few fundraising parties. Heidi (one of the clerks) and I got the idea when we were told we had to paint the stairwell up to Vimy. We both must have been in one of those moods for immediate change because we got the electricians and Boss Beaver together to find out if renovations in the house were possible- leave it to the young females to get these things done! haha. But now with new eager live-ins, plans are moving along quickly and I'm sure that will take up much of our attention in the off hours. 
 
 Other than that I'm looking forward to the arrival of Tim Christie this week who will be replacing me during February. Time is flying up here and it's hard for me to believe that I've just past my half way mark and that I'll be home in just over a week (hopefully!). But saying that, there is a definite need for moisture and sunlight in my life, so I'm pretty anxious to get on that HERC! Also, I need to get away from all the chocolate that is still lying around here from the holidays. John Kivisto keeps bringing more into our office every few days - I think he's trying to fatten us up so that the wolves eyeball us instead! Anyway, that's it for now,
 
Talk to everyone later,

Maria(aka slave driver) - by the way, Frank wrote that! Seems all too similar to what Chad called me! I just tell them that I'm trying to help Environment Canada get their money worth by working the students to the ground! haha.

Maria Vavro
Alert Baseline Operator


It's been a great first week in Alert!  After becoming accustomed to the cold weather and the extreme dryness, everything is smooth sailing as we speak.  I've been "mugged in" to both Vimy and HOC so far awhile meeting great people along the way.  I've managed to produce a full beard and a bald head in the first week here as well.  Maria has been showing me all the requirements of the assistant operator job and things are going well.  My first week at the GAW lab has been very successful and the level of instrumentation is very impressive.  We made a couple of trips to the SST to drop off some of my equipment in preparation for my study.  I'm looking forward to the arrival of Tim and John so I can get my study underway.  That's all for now and I'll talk to you next week.

Frank

Frank Cobbett
Alert Assistant Baseline Operator